VECTOR CORRELATION IN THE PHOTODISSOCIATION OF METAL NITROSYLS

\underline{JEFFREY A. BARTZ}, AMBER L. PEDEN, and RYAN D. KIEDA, Department of Chemistry, Kalamazoo College, 1200 Academy Street, Kalamazoo, MI 49006.

The vector correlation in the photodissociation of metal nitrosyls has been determined using linearly-polarized laser light and velocity-mapped ion imaging. The 225-nm dissociation beam excites a doubly-degenerate metal-to-ligand charge transfer in both η^{5}-C_{5}H_{5}NiNO and Co(CO)_{3}NO. State-resolved detection of the NO product through the A (v’=0) ← X (v”=0) transition reveals that both molecules dissociate promptly with a high degree of vector correlation.